# Coded Downlink Massive Random Access and a Finite de Finetti Theorem

### v1

@@ Abstract
This paper considers a massive connectivity setting in which a base-station (BS) aims to communicate sources $(X_{1},\cdots,X_{k})$ to a randomly activated subset of $k$ users, among a large pool of $n$ users, via a common downlink message. Although the identities of the $k$ active users are assumed to be known at the BS, each active user only knows whether itself is active and does not know the identities of the other active users. A naive coding strategy is to transmit the sources alongside the identities of the users for which the source information is intended, which would require $H(X_{1},\cdots,X_{k})+k\log(n)$ bits, because the cost of specifying the identity of a user is $\log(n)$ bits. For large $n$, this overhead can be significant. This paper shows that it is possible to develop coding techniques that eliminate the dependency of the overhead on $n$, if the source distribution follows certain symmetry. Specifically, if the source distribution is independent and identically distributed (i.i.d.) then the overhead can be reduced to at most $O(\log(k))$ bits, and in case of uniform i.i.d. sources, the overhead can be further reduced to $O(1)$ bits. For sources that follow a more general exchangeable distribution, the overhead is at most $O(k)$ bits, and in case of finite-alphabet exchangeable sources, the overhead can be further reduced to $O(\log(k))$ bits. The downlink massive random access problem is closely connected to the study of finite exchangeable sequences. The proposed coding strategy allows bounds on the relative entropy distance between finite exchangeable distributions and i.i.d. mixture distributions to be developed, and gives a new relative entropy version of the finite de Finetti theorem which is scaling optimal.
@@ I Introduction
I-A Problem Setting
This paper considers the problem of information transmission from a central base-station (BS) to a random subset of $k$ users in a downlink massive random access scenario with a large pool of $n$ users. The users are assumed to be sporadically activated, so that at any given time, only $k\ll n$ users are actively listening to the BS. The BS knows the identities of the active users and wishes to communicate sources $(X_{1},\cdots,X_{k})$ to these active users via a common downlink message; but each user only knows whether it is active itself and does not know which other users are active. We ask the question: What is the minimum rate of downlink common message that allows each active user $i$ to learn the source $X_{i}$ intended for it?
The aforementioned source coding problem, which we refer to as the coded downlink massive random access problem, arises naturally in the context of machine-type communications or Internet-of-Things (IoT), particularly in scenarios where the users are randomly activated and the BS needs to send information (such as the acknowledgement or control messages) to the active users in the downlink [3, 4]. A typical communications protocol in this context consists of two phases. In the first phase, the active users transmit pilots to the BS, which performs activity detection to learn the identities of the active users. In the second phase of the protocol, the BS transmits the information sources to each of the active users via a common downlink broadcast message. This paper is concerned with the fundamental limit of communications in the second phase.
The information sources intended for the set of active users are not necessarily independent (e.g., there may be a cap on the maximum number of users that can be positively acknowledged, or the message may convey the assignment of non-colliding scheduling slots to the active users), but the source distribution typically follows certain symmetry. In this paper, we assume that the user activity patterns are random and symmetric across the pool of all potential users, i.e., no subset of users is preferred over any other subset. Moreover, we assume that the distribution of the information sources intended for the active users is independent of the user activity pattern; further, the source distribution follows a symmetry captured by the notion of exchangeability.
Formally, $(X_{1},\cdots,X_{k})$ is exchangeable if its distribution $p(x_{1},\dots,x_{k})$ is invariant under permutations, i.e. for any bijection $\sigma$ from $\{1,2,\cdots,k\}$ to itself, we have
@@ I-A Problem Setting
Classical examples of exchangeable distributions include independently and identically distributed (i.i.d.) random variables, i.i.d. mixture random variables, and the urn distributions (i.e., sampling with or without replacement from a population).
@@
The downlink massive random access problem would have been trivial if every active user knew the identities of the entire subset of active users. In this case, the $n$ users can each be pre-assigned an index. The BS can then list the source information for the $k$ active users according to the order in which they appear in the index. Such a common message requires only $H(X_{1},\cdots,X_{k})$ bits, but relies on each user knowing the identities of all active users.
@@
The problem becomes much more interesting under the practical scenario in which every active user only knows whether it is active itself and does not know who else is active. In this case, it would appear that the BS needs to send not only $H(X_{1},\cdots,X_{k})$ bits for the source, but also a header of $\log(n)$ bits per active user to describe which user the source is intended for, thus resulting in a $k\log(n)$ bit overhead. Such an overhead can be significant if $n$ is large (e.g., $n=10^{6}$ with $k=10^{3}$ would require $k\log(n)\approx 20$ kbs), especially when the payload for each user is small in comparison. The main insight of this paper is that if the source is exchangeable, it is possible to do much better. In fact, it is possible to develop a coded downlink massive random access strategy at common message rates that do not depend on $n$.
@@
The intuition that makes this possible is that the naive scheme broadcasts too much redundant information. Specifically, each active user needs to recover only its own designated source, and does not care which other users are active, or the information intended for the other users. The naive scheme broadcasts the source information intended for everyone to everyone else. In contrast, the coding strategy developed in this paper reduces the common message rate by taking advantage of the fact that each of the $k$ active users is only interested in the source pertaining to itself, so that each codeword can cover many different instances of the activity patterns and the associated sources.
@@
This paper proposes a coding strategy using a codebook comprised of length-$n$ codewords of different possible realizations of the sources. Each of the $n$ users has a corresponding unique location in the codewords, so that when the identities of the $k$ activity users and their associated sources are revealed, the BS only needs to search over the codebook and broadcast the index of the first codeword that matches the actual sources intended for the $k$ active users to allow them to recover their respective source information.
@@
The main result of this paper is that when the source is distributed according to an exchangeable distribution, it is possible to design such a codebook so that the entropy of the matching index is close to $H(X_{1},\cdots,X_{k})$ with an overhead independent of $n$.
@@
This paper further points out that the problem of constructing a good codebook for an exchangeable source is closely related to the de Finetti theorem and the study of finite exchangeable sequences. This connection arises due to the fact that the codebook construction needs to be symmetric across all subsets of users. Therefore, the distribution used for codebook construction is restricted to be an i.i.d. mixture. The source however can be any exchangeable distribution. Therefore, the problem of constructing a good codebook for an exchangeable source reduces to finding an i.i.d. mixture distribution that is close to an exchangeable distribution in terms of relative entropy. In this paper, we propose a codebook construction method for exchangeable sources and utilize it to prove a relative entropy version of the finite de Finetti theorem that matches the optimal scaling of the classical result [5].
I-B Prior Work
The study of coded downlink for specific sources has been investigated in previous works [6] and [1], where the problem of scheduling active users into non-colliding transmission slots and the problem of assigning active users into categories of fixed sizes are considered respectively. The analyses of code rates in [6], [1] are specific to their respective problems and do not immediately generalize to other source distributions, and in particular do not apply to i.i.d. distributions. In contrast, this paper utilizes new coding techniques to provide results for exchangeable sources in general. Note that both [6] and [1] are examples of exchangeable sources.
@@ I-B Prior Work
It is worth noting that the scheduling code from [6] in conjunction with usual source coding already allow a coding strategy for exchangeable sources at common message rates independent of $n$. The idea is that we can first schedule the $k$ active users into $k$ distinct slots using at most $(k+1)\log(e)$ bits using the strategy in [6], then list the sources information in the order of the slot assignments, so that each active user can receive its correct information. Since the source is exchangeable, the distribution is not affected by the reordering. Therefore a rate of $H(X_{1},\cdots,X_{k})+(k+1)\log(e)$ bits is achievable. This strategy achieves the same $O(k)$ overhead as the coding scheme proposed in this paper for general exchangeable sources, but in many cases, the $O(k)$ overhead can be improved. For example, if the source is i.i.d. or i.i.d. mixture, the coding strategy proposed in this paper achieves an overhead of $O(\log(k))$ bits.
@@
Finite de Finetti theorems have a long history [7, 5, 8]. Let $(X_{1},\cdots,X_{k})$ be an exchangeable sequence with distribution $p(\mathbf{x})$ which takes values in $\mathcal{X}^{k}$ and let $\mathcal{Q}$ be the set of all i.i.d. mixture distributions on $\mathcal{X}^{k}$. We say that the source is $d$-extendable if it has the same distribution as the first $k$ elements of a longer exchangeable sequence $(X_{1},\cdots,X_{k},\cdots,X_{d})$. The classical result of [5] is that
$\min_{q\in\mathcal{Q}}\mathrm{TV}(p,q)\leq\min\left\{\frac{2|\mathcal{X}|k}{d},\frac{k(k-1)}{d}\right\},$ (2)
@@
where TV$(p,q)=2\sup_{\mathcal{X}^{k}}|p(\mathbf{x})-q(\mathbf{x})|$ is the total variation distance. Therefore for any fixed $k$, the total variation distance tends to 00 as $d\to\infty$. In fact, the total variation can go to 00 even if $k,d\to\infty$, as long as $\frac{k}{d}\to 0$ for finite $|\mathcal{X}|$, or $\frac{k^{2}}{d}\to 0$. It is shown in [5] that this scaling is essentially optimal.
@@
Recently, there has been a renewed interest in finding similar bounds as (2) but in terms of relative entropy. In particular, [9, 10, 11] use information-theoretic tools to establish various upper bounds, the strongest of which appears in [11] and shows that
$\min_{q\in\mathcal{Q}}D(p\|q)\leq\frac{k(k-1)}{2(d-k-1)}\log|\mathcal{X}|.$ (3)
However, this result holds only in the finite alphabet setting and has a suboptimal scaling compared to (2). This paper aims to close this gap by proving upper bounds in terms of the relative entropy that match the optimal scaling in (2).
I-C Main Contributions

### v4

@@ Abstract
This paper considers a massive connectivity setting in which a base-station (BS) aims to communicate sources $(X_{1},\cdots,X_{k})$ to a randomly activated subset of $k$ users, among a large pool of $n$ users, via a common message in the downlink. Although the identities of the $k$ active users are assumed to be known at the BS, each active user only knows whether itself is active and does not know the identities of the other active users. A naive coding strategy is to transmit the sources alongside the identities of the users for which the source information is intended. This requires $H(X_{1},\cdots,X_{k})+k\log(n)$ bits, because the cost of specifying the identity of one out of $n$ users is $\log(n)$ bits. For large $n$, this overhead can be significant. This paper shows that it is possible to develop coding techniques that eliminate the dependency of the overhead on $n$, if the source distribution follows certain symmetry. Specifically, if the source distribution is independently and identically distributed (i.i.d.) then the overhead can be reduced to at most $O(\log(k))$ bits, and in case of uniform i.i.d. sources, the overhead can be further reduced to $O(1)$ bits. For sources that follow a more general exchangeable distribution, the overhead is at most $O(k)$ bits, and in case of finite-alphabet exchangeable sources, the overhead can be further reduced to $O(\log(k))$ bits. The downlink massive random access problem is closely connected to the study of finite exchangeable sequences. The proposed coding strategy allows bounds on the Kullback-Leibler (KL) divergence between finite exchangeable distributions and i.i.d. mixture distributions to be developed and gives a new KL divergence version of the finite de Finetti theorem, which is scaling optimal.
@@ I Introduction
This paper considers the problem of information transmission from a central base-station (BS) to a random subset of $k$ users in a downlink massive random access scenario with a large pool of $n$ users. The users are assumed to be sporadically activated, so that at any given time, only $k\ll n$ users are actively listening to the BS. The BS knows the identities of the active users and wishes to communicate sources $(X_{1},\cdots,X_{k})$ to these active users via a downlink common message; but each user only knows whether it is active itself and does not know which other users are active. We ask the question: What is the minimum length of the common message that allows each active user $i$ to learn the source $X_{i}$ intended for it?
I-A Motivation and Problem Setting
The aforementioned source coding problem, which we refer to as the coded downlink massive random access problem, arises naturally in the context of machine-type communications or Internet-of-Things (IoT), particularly in scenarios where the users are randomly activated and the BS needs to send information (such as the acknowledgment or control messages) to the active users in the downlink [3, 4]. A typical communications protocol in this context consists of two phases. In the first phase, the active users transmit pilots to the BS, and the BS performs activity detection to learn the identities of the active users. In the second phase of the protocol, the BS transmits the information sources to each of the active users via a downlink common message. This paper makes the assumption that the user activities are detected correctly in the first phase and focuses on the fundamental limit of communications in the second phase.
The sources intended for the active users are not necessarily independent (e.g., there may be a cap on the maximum number of users that can be positively acknowledged, or the message may convey the assignment of non-colliding scheduling slots to the active users), but under reasonable assumptions, the distribution of the sources typically possesses certain symmetry. In this paper, we assume that the user activity patterns are random and symmetric across the pool of all potential users, i.e., no subset of users is preferred over any other subset. Moreover, we assume that the sources intended for the active users are independent of the user activity pattern. Furthermore, fixing a user activity pattern, we assume that all permutations of any source sequence are equally probable. This symmetry under permutations is known as exchangeability.
Formally, sources $(X_{1},\cdots,X_{k})$ are exchangeable if their distribution $p(x_{1},\dots,x_{k})$ is invariant under permutation, i.e. for any bijection $\sigma$ from $\{1,2,\cdots,k\}$ to itself, we have
@@ I-A Motivation and Problem Setting
Classical examples of exchangeable distributions include independently and identically distributed (i.i.d.) random variables, i.i.d. mixture random variables, and urn distributions (i.e., sampling with or without replacement from a population).
@@
The downlink massive random access problem would have been trivial if every active user knew the identities of the entire subset of active users. In this case, the $n$ users can each be pre-assigned an index. In the downlink phase, the BS can then list the sources for the $k$ active users according to the order in which they appear in the index. Such a common message requires only $H(X_{1},\cdots,X_{k})$ bits, but relies on each user knowing the identities of all active users.
@@
The problem becomes much more interesting under the practical scenario in which every active user only knows whether it is active itself and does not know who else is active. In this case, it would appear that the BS needs to send not only $H(X_{1},\cdots,X_{k})$ bits for the sources, but also a header of $\log(n)$ bits per active user to describe which user each source is intended for, thus resulting in an overhead of $k\log(n)$ bits. Such an overhead can be significant when $n$ is large (e.g., $n=10^{6}$ with $k=10^{3}$ would require $k\log(n)\approx 20$ kilobits), especially if the payload for each user is small in comparison. The main insight of this paper is that if the sources are exchangeable, then it is possible to reduce the overhead significantly as compared to the aforementioned naive scheme. In fact, it is possible to develop a coded downlink massive random access strategy with a common message length that does not depend on $n$.
@@
The intuition that makes this possible is that the naive scheme broadcasts too much redundant information. Specifically, each active user needs to recover only its own designated source, and does not care which other users are active, or the sources intended for the other users. The naive scheme broadcasts the source intended for each of the active users to everyone. In contrast, the coding strategy developed in this paper reduces the common message length by taking advantage of the fact that only the $k$ active users are listening and that each of the $k$ active users is only interested in the source pertaining to itself. This allows a single codeword to cover many different instances of the activity patterns and the associated sources. Equivalently, it allows the common message length needed to convey the $k$ sources to a random subset of $k$ active users to be significantly reduced as compared to the naive scheme.
I-B Proposed Coding Scheme
@@
This paper proposes a coding strategy using a codebook comprised of length-$n$ codewords of different possible realizations of the sources. Each of the $n$ users has a corresponding unique location in the codewords, so that when the identities of the $k$ activity users and their associated sources are revealed, the BS can communicate the sources to the active users by searching over the codebook and broadcasting the index of the first codeword that matches the sources intended for the $k$ active users.
Figure 1: Codebook for transmitting i.i.d. Ber$(\frac{1}{2})$ sources to $k=3$ active users among a total of $n=6$ potential users.
As an example, suppose that we wish to communicate i.i.d. $\mathrm{Ber}(\frac{1}{2})$ sources to $k=3$ active users among a total of $n=6$ potential users. The source entropy is $k$ bits. Instead of using $k+k\log(n)$ bits to send one bit to each of the $k$ active users, we use a codebook comprised of length-$n$ binary vectors, as illustrated for this example in Fig. 1, where each of the $k$ active users has a unique location in the codewords. Suppose that the BS wishes to communicate binary sources $(0,1,1)$ to active users $(1,4,5)$, respectively. The BS would search through the codebook as shown in Fig. 1, select the codeword $4$, (because it is the codeword with the smallest index that has the values $(0,1,1)$ at the active user locations $(1,4,5)$), and broadcast this index $4$ to allow all the active users to recover the sources intended for them. Note that in this coding scheme, a single codeword can be used to communicate many different source-activity pairs. This is because only the active users are listening and the codeword values at the locations of the non-active users are irrelevant. For example, codeword $4$ would also suffice for communicating sources $(0,0,1)$ to active users $(2,3,4)$.
If the user activity patterns are random and the sources are i.i.d. $\mathrm{Ber}(\frac{1}{2})$, we can construct a random codebook with i.i.d. $\mathrm{Ber}(\frac{1}{2})$ entries. In this case, the index of the first codeword that matches the sources at the active users’ locations follows a geometric distribution with parameter $p=\left(\frac{1}{2}\right)^{k}$. Let random variable $T$ denote this index, we have
$\mathrm{Pr}(T=t)=(1-p)^{t-1}p.$ (2)
The entropy of a geometric random variable is bounded from above as
$H(T)$ $=$ $\frac{-p\log(p)-(1-p)\log(1-p)}{p}$ (3)
$\leq$ $\log\left(\frac{1}{p}\right)+\log(e)$ (4)
$=$ $k+\log(e).$ (5)
Thus by using an optimal prefix-free code for compressing $T$, the expected common message length can be shown to be within $\log(e)+1$ bits of the joint entropy of the sources, which is $k$ bits. Most importantly, this same coding strategy works regardless of $n$. In fact, the entropy of the index does not depend on $n$; this contrasts the naive scheme, which would have required an overhead of $k\log(n)$ bits.
@@
The main result of this paper is that the coding strategy illustrated in Fig. 1 can be generalized to arbitrary exchangeable sources. When the sources are exchangeable, it is possible to design a codebook so that the entropy of the matching index is $H(X_{1},\cdots,X_{k})$ plus an overhead independent of $n$.
I-C Connection with Finite de Finetti Theorem
@@
This paper further points out that the problem of constructing a good codebook for exchangeable sources is closely related to the de Finetti theorem and the study of finite exchangeable sequences. This connection arises due to the fact that the codebook construction needs to be symmetric across all subsets of users. Therefore, it is natural to use an i.i.d. mixture distribution to construct the codebook, i.e., each codeword is generated according to the distribution
[…]
The distribution of the sources however can be any exchangeable distribution. Therefore, the problem of constructing a good codebook for exchangeable sources reduces to finding an i.i.d. mixture distribution which is close to the exchangeable distribution in terms of Kullback-Leibler (KL) divergence. In this paper, we propose a codebook construction method for exchangeable sources and utilize it to prove a KL divergence version of the finite de Finetti theorem which matches the optimal scaling of [5].
I-D Related Work
The coding strategy presented in this paper takes inspiration from the collision-free scheduling codes of [6]. In the problem of collision-free scheduling, the BS aims to assign each of the $k$ active users into one of $k$ transmission slots, such that no two active users are assigned to the same slot. Since the BS is not assigning users to specific slots, this problem does not directly correspond to the coded downlink transmission of information sources. Later in this paper, we discuss an alternative scheduling problem in which BS wishes to assign each active user to a specific transmission slot. This alternative scheduling problem would be an instance of the coded downlink transmission. Although the original problem formulation in [6] is different from the setting of this paper, the code constructions bear similarities. Both leverage the fact that a single codeword can be used to schedule many different sets of active users.
@@ I-D Related Work
It is worth noting that the scheduling code from [6] in conjunction with conventional source coding already allow a coding strategy for exchangeable sources at common message lengths independent of $n$. The idea is that we can first schedule the $k$ active users into $k$ distinct slots using at most $(k+1)\log(e)$ bits using the strategy in [6], then list the sources in the order of the slot assignments, so that each active user receives the correct information. Since the sources are exchangeable, the distribution is not affected by the reordering. Therefore a common message length of $H(X_{1},\cdots,X_{k})+(k+1)\log(e)$ bits is achievable. This strategy achieves the same $O(k)$ overhead as the coding scheme proposed in this paper for general exchangeable sources, but in many cases, the $O(k)$ overhead can be improved. For example, if the sources are i.i.d. or i.i.d. mixture, the coding strategy proposed in this paper achieves an overhead of $O(\log(k))$ bits. For sources with uniform distribution over its alphabet, the overhead can be further reduced to $O(1)$ bits.
The problem setting of this paper, when specialized to transmitting i.i.d. bits, is reminiscent of the concept of identification via channels [7] and specifically the idea of identification plus transmission (IT) code in [8], which aims to send a message to one intended user among a large pool of potential users via $n$ uses of a channel of capacity $C$, so that the intended user receives the message with error probability at most $\lambda_{1}$, while the average successful decoding probabilities at all other non-intended users are upper bounded by $\lambda_{2}$. The key result of [8] is that for any positive $(\lambda_{1},\lambda_{2})$, there exists a fixed-length code for the setting in which the number of messages can be up to $2^{nC}$ and the number of potential users can be up to $2^{2^{nC}}$. In other words, the “addressing” information for identifying the intended user in a pool of up to doubly exponential number of potential users can be embedded into the message for the intended user at essentially no extra cost in rate.
The problem setting of this paper differs from that of [8] in that: (i) we consider the case of $k$ intended recipients rather than a single user; (ii) we enforce zero error for the intended users, rather than allowing some positive error probability $\lambda_{1}$; (iii) we assume that all the other users are not listening, thus we do not consider the error probability $\lambda_{2}$; but the number of potential users can be unbounded, rather than being doubly exponential; (iv) we consider variable-length coding instead of fixed-length codes. We show that the “addressing” overhead in the setting of this paper for transmitting i.i.d. Ber$(\frac{1}{2})$ bits to $k$ users is at most a constant, regardless of the size of the pool of potential users.
The coded downlink information transmission problem for massive random access has been considered in the earlier conference version of this work [1, 2], where a specific example of the categorization problem and its generalization are introduced. In this paper, we expand upon [1, 2] to treat more general sources. Furthermore, we draw a connection between the massive random access problem and the theory of finite exchangeable sources.
@@
Finite de Finetti theorems have a long history [9, 5, 10]. Let $(X_{1},\cdots,X_{k})$ be an exchangeable sequence with distribution $p(\mathbf{x})$ which takes values in $\mathcal{X}^{k}$ and let $\mathcal{Q}$ be the set of all i.i.d. mixture distributions on $\mathcal{X}^{k}$. We say that $(X_{1},\cdots,X_{k})$ is $d$-extendable if it has the same distribution as the first $k$ elements of a longer exchangeable sequence $(X_{1},\cdots,X_{k},\cdots,X_{d})$. From [5, Theorems $3$ and $13$], for all $d$-extendable exchangeable sequences $(X_{1},\cdots,X_{k})$ taking values in $\mathcal{X}$, we have
$\min_{q\in\mathcal{Q}}\mathrm{TV}(p,q)\leq\min\left\{\frac{k(k-1)}{2d},\frac{k|\mathcal{X}|}{d}\right\},$ (7)
@@
where $\mathrm{TV}(p,q)=\frac{1}{2}\sum_{\mathbf{x}\in\mathcal{X}^{k}}|p(\mathbf{x})-q(\mathbf{x})|$ is the total variation (TV) distance [^1] This definition of TV distance differs from the definition used in [5] by a factor of $1/2$.. Therefore for any fixed $k$, the TV distance tends to 00 as $d\to\infty$. In fact, the TV distance can go to 00 even if $k\to\infty$, as long as $\frac{k}{d}\to 0$ for finite $|\mathcal{X}|$, or $\frac{k^{2}}{d}\to 0$. It is shown in [5] that this scaling is optimal.
@@
Recently, there has been an interest in finding similar bounds as (7) but in terms of KL divergence instead of TV distance. In particular, [11, 12, 13, 14] utilize information-theoretic arguments to prove various upper bounds on the KL divergence. It is shown in [13] that
$\min_{q\in\mathcal{Q}}D(p\|q)\leq\frac{k(k-1)}{2(d-k+1)}H(X_{1}),$ (8)
where $D(p\|q)$ is the KL divergence. This bound holds for sources with arbitrary alphabet sizes, but the bound depends on the distribution of the sources. In the finite alphabet setting, [14] utilizes the log-sum inequality along with results of [9] to show that
$\min_{q\in\mathcal{Q}}D(p\|q)\leq(|\mathcal{X}|-1)\frac{k(k-1)}{2(d-1)(d-k+1)}.$ (9)
Although the scalings of (8) and (9) match that of [5], both can be improved upon. This paper improves (8) by removing the dependency on the distribution of the sources. Further, in the case of $d=k$, the bound (8) implies a dependency of $O(k^{2})$, while the result of this paper improves the bound to $O(k)$. Moreover, when the alphabet size is finite, in the case of $d=k$, the bound (9) implies a dependency of $O(|\mathcal{X}|k)$. In this paper, we tighten this upper bound to $O(|\mathcal{X}|\log(k))$ using a method-of-types argument.
I-E Main Contributions